OpenStudy (anonymous):

The length of a rectangle is 10 cm less than it's width. what are the dimesons of the rectangle if it's area is 75 square cm.. The length of the rectangle is _?__ cm The width of the rectangle is _?__cm

mathslover (mathslover):

let the width be x given length is 10 cm less than it's width hence length = x-10

mathslover (mathslover):

correct?

OpenStudy (anonymous):

so the length is going to be x-10?

mathslover (mathslover):

yes

OpenStudy (anonymous):

what about the width?

mathslover (mathslover):

width has been taken to be x

mathslover (mathslover):

multiply x-10 and x equate them to 75

OpenStudy (anonymous):

I'm confused what?

mathslover (mathslover):

\[\large{( x-10 )(x)=75}\]

OpenStudy (anonymous):

so that is my answer for the width?

mathslover (mathslover):

no solve the given equation for x

OpenStudy (anonymous):

-5, 15 ??

mathslover (mathslover):

\[\large{x^2-10x = 75}\] \[\large{x^2-10x-75=0}\] \[\large{x^2+15x-5x-75=0}\]

mathslover (mathslover):

5 , - 15

mathslover (mathslover):

check it once again

OpenStudy (anonymous):

so my width is -5, 15??

mathslover (mathslover):

No your width x = 5 since -15 is not possible for x

OpenStudy (anonymous):

the computer said that was inncorrect?

mathslover (mathslover):

oh wait

mathslover (mathslover):

width = 15

OpenStudy (anonymous):

and the length is x-15 right??

mathslover (mathslover):

length = x- 10 = 15 - 10 = 5

OpenStudy (anonymous):

so my length is 5 and width is 15

OpenStudy (anonymous):

??

mathslover (mathslover):

right